Session: Sep 2000 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Wed Sep 20 11:13:43 METDST 2000) Success
No known problems.

Torun (K.Borkowski at Wed Sep 20 13:06:05 METDST 2000) Success
No known problems.

Noto (C.Trigilio at Wed Sep 20 14:14:01 METDST 2000) Success

Effelsberg (A. Kraus at Thu Sep 21 14:49:24 METDST 2000) Minor failures
LCP and RCP are exchanged! The receivers RCP-channel has much higher system temperature than LCP. More details on the 5cm-session page.

Onsala (A. Polatidis at Mon Sep 25 00:48:20 METDST 2000) Success
REcording LINEAR POLARIZATION at both IF1 and IF2

Jodrell_Bank (Alastair Gunn at Tue Sep 26 06:29:36 METDST 2000) Minor failures
Stopped schedule soon after start to solve previous problems with receiver and FS (see experiment EM039). FS restarted and schedule run from 10:03 UT. Data should be permanently good after this time.